Solve the equation
3/2(-6x+8)=-1

Answer by stanbon(75887)

Solve the equation
3/2(-6x+8)=-1
(-6x+8) = 2/3
-6x = (2/3)-(24/3)
-6x = -22/3
x = 22/18
x = 11/9
--------------------------------
Cheers,
Stan H.